In a multiple regression with four explanatory variables and 111 observations, it is found that SSR = 4.85 and SST = 7.89.

a. Calculate the standard error of the estimate se. (Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

b. Calculate the coefficient of determination R2. (Round your answer to 4 decimal places.)

c. Calculate adjusted coefficient of determination R2. (Round your answer to 4 decimal places.)
Answer #1

a) SSE = SST - SSR = 7.89 - 4.85 = 3.04

The standard error of the estimate =

b) The coefficient of determination =

c) The adjusted coefficient of determination =